OpenStudy (anonymous):

Solve square root of x plus 1 end root plus 5 equals x

OpenStudy (anonymous):

Could you express it in terms of numbers?

OpenStudy (anonymous):

\[\sqrt{x+1} + 5=x\]

OpenStudy (anonymous):

Try substracting by five, then taking x-5 to the power of 2, to get rid of the radical.

OpenStudy (anonymous):

Subtracting*

OpenStudy (comm.dan):

No solution because you can't have a plus and minus square root for an imperfect square. @JessieJakeway

OpenStudy (anonymous):

Thank you! @Comm.Dan You're the best! :)

OpenStudy (comm.dan):

No problem!! Any more questions?? @JessieJakeway

OpenStudy (anonymous):

Simplify the square root of 128 end root plus 2 square root of 112 end root minus 2 square root of 98 end root minus 6 square root of 28

OpenStudy (comm.dan):

Could u please put it into terms of numbers?

OpenStudy (anonymous):

\[\sqrt{128}+2\sqrt{112}-2\sqrt{98}-6\sqrt{28}\]

OpenStudy (comm.dan):

Thx

jimthompson5910 (jim_thompson5910):

x = 8 is a solution to \[\Large \sqrt{x+1} + 5=x\] since plugging in x = 8 gives you a true equation \[\Large \sqrt{x+1} + 5=x\] \[\Large \sqrt{8+1} + 5=8\] \[\Large \sqrt{9} + 5=8\] \[\Large 3 + 5=8\] \[\Large 8=8\] which is true, so that verifies that x = 8 is a solution

OpenStudy (comm.dan):

Where did u get the 8 from? @jim_thompson5910

OpenStudy (comm.dan):

 @JessieJakeway